P4. An aluminum wire of diameter 1.00 mm is put under a tension of 375 N. If a 440. Hz driving force sets up a wave that propagates along this wire, calculate the wavelength.
Answer #1

speed of a wave on a stretched string :

v = sqrt(T/(m/L))

density of aluminium, d = 2700 kg/m3

Now, mass , m = d*pi*r^2*L

So, m/L = d*pi*r^2 = 2700*pi*0.001^2 = 8.48*10^-3 kg/m

So, v = sqrt(375/(0.00848))

= 210.3 m/s

So, wavelength = v/f = 210.3/440 = 0.478 m <------ answer